I need help on this question:
A total of $30,000 is invested in two accounts paying 3% and 4% annual intrest. The combined annual interest is $1020
A. Write a system of linear equations that represent the situation.
B. How much of the $30,000 is invested in each account?
Thank you.

Answer by RAY100(1637)

Let x, y = amounts invested
.
(1) x+y = $30,000,,,,,x=30,000-y
.
(2) .03x +.04y =1020
.
substitute (1) into (2)
.
.03(30,000-y) + .04y = 1020
.
900 - .03y +.04y = 1020
.
.01y = 120
.
y= $ 12,000,,,,,,,,x=30,000-12,000= $ 18,000
.
check,,,,.03(18,000) + .04(12,000) = 540 +480 = 1020,,,,,,ok